OpenStudy (candacetoocool):

What is the expression in factored form? x2 + 14x + 48 A. (x + 6)(x - 8) B. (x + 8)(x - 6) C. (x - 8)(x - 6) D. (x + 6)(x + 8)

OpenStudy (raden):

you need 2 numbers whose product is 48 and when these are added get 14

OpenStudy (raden):

find that numbers ?

OpenStudy (anonymous):

D. (x + 6)(x + 8)

OpenStudy (raden):

yes, correct ^

OpenStudy (anonymous):

thx..please click best response if any people answer true...

OpenStudy (unklerhaukus):

\[x^2 + 14x + 48\]\[=(x+a)(x+b)=x^2+ax+bx+ab=x^2+(a+b)x+ab\] \[\qquad\qquad\qquad 14=a+b\] \[\qquad\qquad \qquad ab=48\] all the pairs of factors of \(48\) are ; \((\pm1)(\pm48),~(\pm2)(\pm24),~(\pm3)(\pm16),~(\pm4)(\pm12),~(\pm6)(\pm8)\) \(a ~~\& ~~b\) are the factor pair that add to \(14\)